FIFO Method, Unit Cost, Valuation of Goods Transferred Out and Ending Work in Process

Dama Company produces women's blouses and uses the FIFO method to account for its manufacturing costs. The product Dama makes passes through two processes: Cutting and Sewing. During April, Dama's controller prepared the following equivalent units schedule for the Cutting Department:

Direct Materials Conversion Costs
Units started and completed 40,000 40,000
Units, beginning work in process:
20,000 × 0%
20,000 × 50% 10,000
Units, ending work in process:
20,000 × 100% 20,000
20,000 × 25% 5,000
Equivalent units of output 60,000 55,000

Costs in beginning work in process were direct materials, $25,000; conversion costs, $80,000. Manufacturing costs incurred during April were direct materials, $240,000; conversion costs, $352,000.

Required:

1. Prepare a physical flow schedule for April.

Dama Company
Physical Flow Schedule
Units to account for:
Units, beginning work in process
Units started
Total units to account for
Units accounted for:
Started and completed
Units, beginning work in process
Units, ending work in process
Total units accounted for

Feedback

1. A physical flow schedule shows the units to account for and what happened to them.

2. Compute the cost per equivalent unit for April. If required, round your answer to the nearest cent.
$ per equivalent unit

3. Determine the cost of ending work in process and the cost of goods transferred out.

Cost of ending work in process $
Cost of goods transferred out $

Feedback

2. The equivalent units schedule measures the output of the period. A fully completed unit is counted as a unit of output. Output for a unit in ending work in process is counted by its degree of completion.

3. Unit cost = Costs of the period ÷ Output of the period. The cost of goods (services) transferred out is the unit cost multiplied by the units completed. The cost of EWIP is the unit cost multiplied by the equivalent units found in EWIP.

4. Prepare the journal entry that transfers the costs from Cutting to Sewing.

Work in Process-Sewing
Work in Process-Cutting
Answer #1

Solution Dama company physical Flow schedule process 20000 units to account for : units, beginning work in units started Tota27 cost per equivalent unit of Material = 240000/60000 = 4 per equivalent unit Cost per equivalent unit of conversion = 352003) Cost of ending work in process = [20000 *4 + 5000X 6-4] = 12000 Cost of goods Transfered out = [ 40000 x4 + 50000 X6-4]+- 4). Journal entry accounts and explation Debit 1 Credit 585000 work work in in process process Sewing cutting 585000